OpenStudy (anonymous):

Fan and medal? Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (haseeb96):

C is the correct answer .

OpenStudy (anonymous):

Thankssss

OpenStudy (haseeb96):

Just multiply the first equation by 2 then it will be 10x+2y=18 3x+2y=4 Now subtract both the equations then u will get 7x. =14 x=14/7 x=2 Now put x=2 into equation#1 and solve it then u will get y=-1 So the correct answer will be(x,y)=(2,-1)